OpenStudy (anonymous):

The product of 9 and a number is 12 less than three times that number. What is the number?

OpenStudy (anonymous):

@Secret-Ninja @dtan5457

OpenStudy (secret-ninja):

So... \[9 \times x = 12<3x\]

OpenStudy (anonymous):

9x=4x?

OpenStudy (secret-ninja):

You just have to get the x on its own side and solve.

OpenStudy (dtan5457):

I remember.

OpenStudy (dtan5457):

My equation would be 9x-12=3x

OpenStudy (secret-ninja):

Oh darn I got to go. Bye guys >.<

OpenStudy (anonymous):

ok bye ;/

OpenStudy (dtan5457):

I'll take it from here

OpenStudy (dtan5457):

;p

OpenStudy (dtan5457):

9x-12=3x solve for x

OpenStudy (anonymous):

x=-6

OpenStudy (anonymous):

?

OpenStudy (dtan5457):

Uhh no. 9x-12=3x isolating x gives you 6x=12

OpenStudy (dtan5457):

x=?

OpenStudy (dtan5457):

remember when i said x has to be x, not 6x, -x, etc?

OpenStudy (anonymous):

? how

OpenStudy (dtan5457):

6x=12 if you want x to be alone, you'll divide everything by 6.

OpenStudy (anonymous):

so x would =6

OpenStudy (dtan5457):

im pretty sure 12 divided by 6 is not 6 ;p

OpenStudy (anonymous):

oh lol it would be 2

OpenStudy (dtan5457):

Yep. and you can even check "The product of 9 and a number is 12 less than three times that number." 9 times 2=18 2 times =6 18 is 12 more than 6 correct :)

OpenStudy (anonymous):

so the number would be 2 right

OpenStudy (anonymous):

ok im going to turn in my test

OpenStudy (dtan5457):

yes.

OpenStudy (anonymous):

thanks guys i Got an 80%

OpenStudy (dtan5457):

wow just 80..from the questions i answered??

OpenStudy (anonymous):

yes thats

OpenStudy (anonymous):

*thanks

OpenStudy (dtan5457):

i dont see what answers i did wrong but if your ok with it...that's cool

OpenStudy (anonymous):

no it was not you it was me the only question that i answer my self i had five question i only missed 1 out 5 so i get 4/5

OpenStudy (anonymous):

thanks again

OpenStudy (secret-ninja):

Sorry I left so unexpectedly. Good job you guys, you did great. :D

OpenStudy (anonymous):

You guys were on the right track I believe the answer is -2 because 9*(-2)=-18 and 3*(-2) is -6 and -6 is 12 less than -18. I took this test and I got a 100% using -2 as the answer. I am in the 7th grade too